# A Unified Photometric Redshift Calibration for Weak Lensing Surveys using the Dark Energy Spectroscopic Instrument

### v1

@@ Abstract
1 Introduction
Since the first robust detection in the early 2000s (see, e.g., Bacon et al., 2000; Kaiser et al., 2000; Van Waerbeke et al., 2000; Wittman et al., 2000), weak gravitational lensing by large-scale structure has become a major cosmological probe. By studying the shape perturbations of the images of galaxies, we can obtain detailed measurements of the growth of structure over cosmic time. In the coming decade, experiments such as the Euclid satellite (Euclid Collaboration et al., 2025), the Nancy Grace Roman Space Telescope (Akeson et al., 2019), and the Vera C. Rubin Observatory (The LSST Dark Energy Science Collaboration et al., 2018) will image hundreds of millions of galaxies and detect weak gravitational lensing with unprecedented statistical precision. The increasing fidelity of these experiments puts stringent requirements on potential systematic errors, especially the calibration of photometric redshifts.
@@ 1 Introduction
For stage-III experiments such as the Dark Energy Survey (DES; Flaugher et al., 2015; Abbott et al., 2022), the Subaru Hyper Suprime-Cam Survey (HSC; Aihara et al., 2018; More et al., 2023), and the Kilo-Degree Survey (KiDS; Kuijken et al., 2015; Asgari et al., 2021) the redshift distribution in each bin must be characterized to a precision of $\sigma_{\bar{z}}/(1+\bar{z})\sim 0.01$. For future stage-IV experiments such as the year-10 weak lensing analysis of the Vera C. Rubin Observatory, the requirement are as stringent as $\sigma_{\bar{z}}/(1+\bar{z})=0.001$ (The LSST Dark Energy Science Collaboration et al., 2018).
@@
Different methods for the calibration of photometric redshifts have been developed (see Newman & Gruen, 2022, and references therein). One approach is to use high-quality, ideally spectroscopic, redshifts for a subset of weak lensing galaxies to determine $n(z)$ “directly.” A common difficulty with this so-called direct calibration is that the sample of galaxies with high-quality redshifts is typically not representative of the weak lensing sample. This requires additional steps to mitigate this bias, with self-organizing maps (SOMs, Kohonen, 2001) being a commonly used tool (Masters et al., 2019; Buchs et al., 2019; Wright et al., 2020). A SOM is an unsupervised machine-learning technique for dimensionality reduction. In the context of photometric redshifts, it is typically used to map the high-dimensional color space of galaxies into a two-dimensional representation whereby galaxies belonging to the same or nearby SOM cells have similar colors. This allows for the correction of non-representative samples of galaxies by comparing the distribution of SOM cells in the calibration sample to the wider weak lensing sample. A second widely-used approach is known as clustering redshifts. Most commonly, one uses angular cross-correlations between the weak lensing source galaxies and samples of objects with spectroscopic redshifts. The measured cross-correlation is proportional to $n(z)b(z)$, where $b(z)$ is the large-scale bias of the weak lensing sample. Assuming a reasonably flexible model for $b(z)$, one can then estimate $n(z)$.
@@
The Dark Energy Spectroscopic Instrument (DESI; Levi et al., 2013; DESI Collaboration et al., 2016a, b, 2022; Silber et al., 2023; Miller et al., 2024; Poppett et al., 2024; DESI Collaboration et al., 2024a, b) is a stage-IV spectroscopic cosmology survey using a highly multiplexed fiber-fed spectrograph on the Mayall Telescope. DESI is able to measure spectroscopic redshifts for around $5000$ galaxies at a time (Guy et al., 2023; Schlafly et al., 2023). DESI, thus, is ideally poised to substantially improve photometric redshift calibration via both the direct and clustering redshift method. Recently, Ratajczak et al. (2025) conducted a custom evaluation of all spectra obtained by DESI within two independent regions, one that encompasses the COSMOS field, and one that encompasses the XMM-LSS field. The targets for DESI spectroscopy in these two areas were selected for a broader range of analyses than the core large-scale structure samples, and thus represent a more representative sampling of the full galaxy population than the spectra over the full DESI footprint. At $r>16$, the surface density is $208\%$ higher than that for the DESI main sample for relatively red galaxies with gr>1g-r>1, and $250\%$ larger for relatively blue galaxies with color gr<1g-r<1. The sample of extended objects in an 7 $\mathrm{deg}^{2}$ annulus centered on the COSMOS field is roughly $24\%$ complete down to $i<23$.
@@
In this work, we use the new DESI COSMOS-XMM catalog (Ratajczak et al., 2025) to perform an independent, unified photometric redshift calibration of the three leading stage-III weak lensing surveys, DES, HSC, and KiDS. A companion paper by Blanco et al. (in prep.) studies in more detail the potential for DESI to directly calibrate the KiDS data set. We show that DESI significantly increases the number of spectroscopic redshifts available in the COSMOS field. At the same time, DESI data does not fill the entire color space and misses redshifts above $z\sim 1.6$. In this paper, while we use the direct calibration method, we also introduce and verify a new technique that is based on importance weights instead of SOMs. Additionally, we compare our results against the fiducial calibration results obtained from the different weak lensing surveys. Our analysis method is substantially different from previous work and, additionally, we use a separate new catalog of spectroscopic redshifts from DESI, independent of the redshift compilations used in previous works. Thereby, our study addresses recent cosmological tensions from weak lensing surveys, most notably the $S_{8}$-tension (see Abdalla et al., 2022, and references therein). Finally, in a separate study, Ruggeri et al. (in prep.), we instead perform redshift calibrations with clustering redshifts using the DESI large-scale structure catalogs.
@@
This paper is organized as follows. Section 2 briefly describes the different data sets used. In section 3, we describe our calibration method which is verified in section 4. Our main results, estimates of the redshift distributions $n(z)$ for DES, HSC, and KiDS, are presented in section 5. We will discuss our results in section 6 before concluding in section 7. Throughout this work, we will assume cosmological parameters consistent with the Planck Collaboration et al. (2020) analysis of cosmic microwave background.
2 Data
@@ 2.1 Calibration Data
Our calibration catalog is based on three distinct catalogs: the COSMOS2020 photometric catalog (Weaver et al., 2022), the deep photometric COSMOS catalog of the DESI Legacy Survey DR10 (Zhou et al., in prep.), and the spectroscopic DESI COSMOS-XMM catalog (Ratajczak et al., 2025). The calibration catalog uses the COSMOS2020 catalog as a basis, adding in Legacy Survey and DESI information for individual objects where available. The deep Legacy Survey catalog in COSMOS has $5\sigma$ imaging depths of $26.4$, $26.0$, $25.7$, and $24.9$ for the grizgriz bands. Thus, both COSMOS2020 and the Legacy Survey catalog are significantly deeper than the weak lensing catalogs considered here. The DESI COSMOS-XMM redshift catalogs include 307,934 unique spectra in total over the two $16$ deg2 fields. This publicly available catalog includes redshift classifications and matches to the DESI Legacy Survey DR10 (Dey et al., 2019), HSC PDR3 (Aihara et al., 2022), and COSMOS2020 (Weaver et al., 2022) photometry.
@@
From COSMOS2020, we select all objects in the $2\,\mathrm{deg}^{2}$ COSMOS area (FLAG_COMBINED==0) with valid redshifts (lp_type!=-9). We use the LePhare photometric redshifts, i.e., lp_zBEST, for galaxies and stars and lp_zq for X-ray sources. From the DESI COSMOS-XMM catalog, we keep all targets but consider a spectroscopic redshift unsuccessful if the redshift is not secure (NOT QUALITY_Z) (Ratajczak et al., 2025). We associate sources in the COSMOS2020 catalog with the DESI Legacy Survey DR10 catalog and the DESI catalog based on a $1\,\mathrm{arcsec}$ matching radius. In case multiple exposures are present in the DESI catalog within a given matching radius, we give preference to successful redshifts in DESI’s large-scale structure target classes, i.e., BGS, LRG, ELG, and QSO, and then other successful redshifts. Ultimately, each object in our calibration catalog will have COSMOS2020 photometry and photometric redshifts, a subset has Legacy Survey photometry, and an even smaller subset of that will have spectra, and possibly redshifts, from DESI.
@@
In our analysis, we will use a subset of that calibration sample with the goal of extracting as much redshift information from the secure DESI spectroscopic redshifts. The basic idea is that galaxy populations with sufficient coverage by DESI will exclusively be calibrated with DESI redshifts. In Fig. 2, we display basic properties of the calibration sample such as redshift and magnitude distributions. We also demonstrate the weighting procedure that is described in more detail in section 3.1.
2.2 Dark Energy Survey
@@
The year 3 data release of the Dark Energy Survey (DES Y3) targets $\sim 4000\,\mathrm{deg}^{2}$ on the sky with an effective number density of $5.6\,\mathrm{arcmin}^{-2}$ (Gatti et al., 2021). Galaxies are assigned into $4$ tomographic bins based on a SOM of the $3$-band rizriz photometry (Myles et al., 2021). In this work, we utilize the so-called Balrog catalogs (Everett et al., 2022) in which objects from deep photometric fields, such as COSMOS, are artificially injected into the DES Y3 footprint, including photometric noise. The artificial injections are processed with the same pipeline as the real data, allowing us to estimate the probability of each galaxy in the COSMOS field to pass the DES Y3 selection cuts and be assigned a certain tomographic bin as well as their lensing weight. Finally, the fiducial redshift calibration is based on a combination of SOM-based direct and clustering redshift calibration (Myles et al., 2021) and also accounts for the effects of blending (MacCrann et al., 2022).
2.3 Hyper Suprime-Cam
@@
The Hyper Suprime-Cam first-year data (HSC Y1), based on the intermediate data release S16a, covers $137\,\mathrm{deg}^{2}$ with an effective number density of $\sim 22\,\mathrm{arcmin}^{-2}$ (Mandelbaum et al., 2018). Galaxies are assigned into $4$ tomographic bins based on photometric Ephor AB redshifts from $5$-band grizygrizy photometry (Hikage et al., 2019). For our analysis, we make use of the S17A Wide-COSMOS catalog of the COSMOS field that is based on observations at the depth of the full survey, contains all objects passing the fiducial weak lensing selections cuts, and is matched against the COSMOS2015 catalog (Weaver et al., 2022). In the default HSC first-year photometric redshift calibration, targets are weighted by the product of their lensing weights and a SOM-based weight to partially correct for cosmic variance in the COSMOS field. The fiducial $n(z)$ is then based on the weighted histogram of COSMOS2015 multi-band photometric redshifts, i.e., no spectroscopic redshifts were employed (Hikage et al., 2019).
2.4 Kilo-Degree Survey
@@
The KiDS-1000 data release of the Kilo-Degree Survey covers $\sim 1000\,\mathrm{deg}^{2}$ with an effective number density of $6.2\,\mathrm{arcmin}^{-2}$ (Giblin et al., 2021). A total of $5$ tomographic bins are assigned based on BPZ photometric redshifts from the $ugriZYJHK_{s}$ optical to near-infrared bands. For our analysis, we use a photometric catalog of the COSMOS field at a higher depth than the main survey that has been artificially degraded to simulate the depth of the KiDS-1000 weak lensing catalog (A. Wright, personal communication). Unlike for DES Y3 and HSC Y1, this KiDS-1000 COSMOS catalog has not been processed with the weak lensing pipeline. To estimate the impact of weak lensing cuts and weights, we also make use of the DR 4.1 photometric catalogs in combination with the KiDS-1000 weak lensing SOM-gold catalog, as described in section 3.2. Finally, the fiducial photometric redshift calibration has been described in Hildebrandt et al. (2021) and is based on a combination of SOM and clustering redshifts.
3 Methods
@@ 3.1 Incompleteness Correction
3.1.1 Fiber Assignment Probabilities
Ideally, the fiber assignment probability would be computed from first principles if the targeting criteria were well-defined and we had random realizations of the fiber assignment such as for the DESI large-scale structure catalogs (Lasker et al., 2025). In practice, the DESI COSMOS-XMM catalog contains targets from a large heterogeneous set of secondary target programs, often without alternative realizations of fiber assignments.
@@ 3.1.1 Fiber Assignment Probabilities
The values predicted by the trained network can be interpreted as the fiber fraction $p_{\mathrm{f}}$ given an object’s coordinates and photometry. To avoid spurious overfitting we employ a K-fold approach by always training the networks on $90\%$ of the data and calculating $p_{\mathrm{f}}$ for the remaining $10\%$, i.e., $K=10$. Finally, this process is repeated $100$ times for the whole sample such that each individual galaxy has $100$ predictions. We typically find the scatter between different neural network predictions $\sigma_{p_{\mathrm{f}}}$ for the same galaxy to be of order $5\%$ or less. Thus, after averaging those predictions, the scatter is of order $0.5\%$ or less. We note that while the network has access to the coordinates of objects, it is unlikely to faithfully reproduce small-scale fiber assignment effects such as fiber collisions and the influence of higher-priority targets (Lasker et al., 2025). This is not a concern for our work since we do not study small-scale clustering. Instead, we are mainly concerned with having accurate mean fiber probabilities as a function of photometry within overlap regions with weak lensing surveys. Finally, we observe that the raw $p_{\mathrm{f}}$ do not reproduce the mean fiber rate $f_{\mathrm{f}}$ in $5\%$ bins of $p_{\mathrm{f}}$ perfectly and instead show small offsets. We mitigate this by manually adjusting the raw $p_{\mathrm{f}}$ values via $p_{\mathrm{f}}\rightarrow f_{\mathrm{f}}(p_{\mathrm{f}})$.
3.1.2 Importance Weights

### v2

@@ Abstract
I Introduction
Since the first robust detection in the early 2000s (see, e.g., Bacon2000_MNRAS_318_625; Kaiser2000_astro_ph_3338; VanWaerbeke2000_AA_358_30; Wittman2000_Natur_405_143), weak gravitational lensing by large-scale structure has become a major cosmological probe. By studying the shape perturbations of the images of galaxies, we can obtain detailed measurements of the growth of structure over cosmic time. In the coming decade, experiments such as the Euclid satellite (EuclidCollaboration2025_AA_697_1), the Nancy Grace Roman Space Telescope (Akeson2019_arXiv_1902_5569), and the Vera C. Rubin Observatory (TheLSSTDarkEnergyScienceCollaboration2018_arXiv_1809_1669) will image hundreds of millions of galaxies and detect weak gravitational lensing with unprecedented statistical precision. The increasing fidelity of these experiments puts stringent requirements on potential systematic errors, especially the calibration of photometric redshifts.
@@ I Introduction
For stage-III experiments such as the Dark Energy Survey (DES; Flaugher2015_AJ_150_150; Abbott2022_PhRvD_105_3520), the Subaru Hyper Suprime-Cam Survey (HSC; Aihara2018_PASJ_70_4; More2023_PhRvD_108_3520), and the Kilo-Degree Survey (KiDS; Kuijken2015_MNRAS_454_3500; Asgari2021_AA_645_104) the redshift distribution in each bin must be characterized to a precision of $\sigma_{\bar{z}}/(1+\bar{z})\sim 0.01$. For future stage-IV experiments such as the year-10 weak lensing analysis of the Vera C. Rubin Observatory, the requirement are as stringent as $\sigma_{\bar{z}}/(1+\bar{z})=0.001$ (TheLSSTDarkEnergyScienceCollaboration2018_arXiv_1809_1669).
@@
Different methods for the calibration of photometric redshifts have been developed (see Newman2022_ARAA_60_363, and references therein). One approach is to use high-quality, ideally spectroscopic, redshifts for a subset of weak lensing galaxies to determine $n(z)$ “directly.” A common difficulty with this so-called direct calibration is that the sample of galaxies with high-quality redshifts is typically not representative of the weak lensing sample. This requires additional steps to mitigate this bias, with self-organizing maps (SOMs, Kohonen2001_som_book) being a commonly used tool (Masters2019_ApJ_877_81; Buchs2019_MNRAS_489_820; Wright2020_AA_637_100). A SOM is an unsupervised machine-learning technique for dimensionality reduction. In the context of photometric redshifts, it is typically used to map the high-dimensional color space of galaxies into a two-dimensional representation whereby galaxies belonging to the same or nearby SOM cells have similar colors. This allows for the correction of non-representative samples of galaxies by comparing the distribution of SOM cells in the calibration sample to the wider weak lensing sample. A second widely-used approach is known as clustering redshifts. Most commonly, one uses angular cross-correlations between the weak lensing source galaxies and samples of objects with spectroscopic redshifts. The measured cross-correlation is proportional to $n(z)b(z)$, where $b(z)$ is the large-scale bias of the weak lensing sample. Assuming a reasonably flexible model for $b(z)$, one can then estimate $n(z)$.
@@
The Dark Energy Spectroscopic Instrument (DESI; Levi2013_arXiv_1308_0847; DESICollaboration2016_arXiv_1611_0036; DESICollaboration2016_arXiv_1611_0037; DESICollaboration2022_AJ_164_207; Silber2023_AJ_165_9; Miller2024_AJ_168_95; Poppett2024_AJ_168_245; DESICollaboration2024_AJ_167_62; DESICollaboration2024_AJ_168_58) is a stage-IV spectroscopic cosmology survey using a highly multiplexed fiber-fed spectrograph on the Mayall Telescope. DESI is able to measure spectroscopic redshifts for around $5000$ galaxies at a time (Guy2023_AJ_165_144; Schlafly2023_AJ_166_259). DESI, thus, is ideally poised to substantially improve photometric redshift calibration via both the direct and clustering redshift method. Recently, Ratajczak2025_arXiv_2508_9286 conducted a custom evaluation of all spectra obtained by DESI within two independent regions, one that encompasses the COSMOS field, and one that encompasses the XMM-LSS field. The targets for DESI spectroscopy in these two areas were selected for a broader range of analyses than the core large-scale structure samples, and thus represent a more representative sampling of the full galaxy population than the spectra over the full DESI footprint. At $r>16$, the surface density is $208\%$ higher than that for the DESI main sample for relatively red galaxies with gr>1g-r>1, and $250\%$ larger for relatively blue galaxies with color gr<1g-r<1. The sample of extended objects in an 7 $\mathrm{deg}^{2}$ annulus centered on the COSMOS field is roughly $24\%$ complete down to $i<23$.
@@
In this work, we use the new DESI COSMOS-XMM catalog (Ratajczak2025_arXiv_2508_9286) to perform an independent, unified photometric redshift calibration of the three leading stage-III weak lensing surveys, DES, HSC, and KiDS. A companion paper by Blanco et al. (in prep.) studies in more detail the potential for DESI to directly calibrate the KiDS data set. We show that DESI significantly increases the number of spectroscopic redshifts available in the COSMOS field. At the same time, DESI data does not fill the entire color space and misses redshifts above $z\sim 1.6$. In this paper, while we use the direct calibration method, we also introduce and verify a new technique that is based on importance weights instead of SOMs. Additionally, we compare our results against the fiducial calibration results obtained from the different weak lensing surveys. Our analysis method is substantially different from previous work and, additionally, we use a separate new catalog of spectroscopic redshifts from DESI, independent of the redshift compilations used in previous works. Thereby, our study addresses recent cosmological tensions from weak lensing surveys, most notably the $S_{8}$-tension (see Abdalla2022_JHEAp_34_49, and references therein). Finally, in a separate study, Ruggeri et al. (in prep.), we instead perform redshift calibrations with clustering redshifts using the DESI large-scale structure catalogs.
@@
This paper is organized as follows. Section II briefly describes the different data sets used. In section III, we describe our calibration method which is verified in section IV. Our main results, estimates of the redshift distributions $n(z)$ for DES, HSC, and KiDS, are presented in section V. We will discuss our results in section VI before concluding in section VII. Throughout this work, we will assume cosmological parameters consistent with the PlanckCollaboration2020_AA_641_6 analysis of cosmic microwave background.
II Data
@@ II.1 Calibration Data
Our calibration catalog is based on three distinct catalogs: the COSMOS2020 photometric catalog (Weaver2022_ApJS_258_11), the deep photometric COSMOS catalog of the DESI Legacy Survey DR10 (Zhou et al., in prep.), and the spectroscopic DESI COSMOS-XMM catalog (Ratajczak2025_arXiv_2508_9286). The calibration catalog uses the COSMOS2020 catalog as a basis, adding in Legacy Survey and DESI information for individual objects where available. The deep Legacy Survey catalog in COSMOS has $5\sigma$ imaging depths of $26.4$, $26.0$, $25.7$, and $24.9$ for the grizgriz bands. Thus, both COSMOS2020 and the Legacy Survey catalog are significantly deeper than the weak lensing catalogs considered here. The DESI COSMOS-XMM redshift catalogs include 307,934 unique spectra in total over the two $16$ deg2 fields. This publicly available catalog includes redshift classifications and matches to the DESI Legacy Survey DR10 (Dey2019_AJ_157_168), HSC PDR3 (Aihara2022_PASJ_74_247), and COSMOS2020 (Weaver2022_ApJS_258_11) photometry.
@@
From COSMOS2020, we select all objects in the $2\,\mathrm{deg}^{2}$ COSMOS area (FLAG_COMBINED==0) with valid redshifts (lp_type!=-9). We use the LePhare photometric redshifts, i.e., lp_zBEST, for galaxies and stars and lp_zq for X-ray sources. From the DESI COSMOS-XMM catalog, we keep all targets but consider a spectroscopic redshift unsuccessful if the redshift is not secure (NOT QUALITY_Z) (Ratajczak2025_arXiv_2508_9286). We associate sources in the COSMOS2020 catalog with the DESI Legacy Survey DR10 catalog and the DESI catalog based on a $1\,\mathrm{arcsec}$ matching radius. In case multiple exposures are present in the DESI catalog within a given matching radius, we give preference to successful redshifts in DESI’s large-scale structure target classes, i.e., BGS, LRG, ELG, and QSO, and then other successful redshifts. Ultimately, each object in our calibration catalog will have COSMOS2020 photometry and photometric redshifts, a subset has Legacy Survey photometry, and an even smaller subset of that will have spectra, and possibly redshifts, from DESI.
@@
In our analysis, we will use a subset of that calibration sample with the goal of extracting as much redshift information from the secure DESI spectroscopic redshifts. The basic idea is that galaxy populations with sufficient coverage by DESI will exclusively be calibrated with DESI redshifts. In Fig. 2, we display basic properties of the calibration sample such as redshift and magnitude distributions. We also demonstrate the weighting procedure that is described in more detail in section III.1.
II.2 Dark Energy Survey
@@
The year 3 data release of the Dark Energy Survey (DES Y3) targets $\sim 4000\,\mathrm{deg}^{2}$ on the sky with an effective number density of $5.6\,\mathrm{arcmin}^{-2}$ (Gatti2021_MNRAS_504_4312). Galaxies are assigned into $4$ tomographic bins based on a SOM of the $3$-band rizriz photometry (Myles2021_MNRAS_505_4249). In this work, we utilize the so-called Balrog catalogs (Everett2022_ApJS_258_15) in which objects from deep photometric fields, such as COSMOS, are artificially injected into the DES Y3 footprint, including photometric noise. The artificial injections are processed with the same pipeline as the real data, allowing us to estimate the probability of each galaxy in the COSMOS field to pass the DES Y3 selection cuts and be assigned a certain tomographic bin as well as their lensing weight. Finally, the fiducial redshift calibration is based on a combination of SOM-based direct and clustering redshift calibration (Myles2021_MNRAS_505_4249) and also accounts for the effects of blending (MacCrann2022_MNRAS_509_3371).
II.3 Hyper Suprime-Cam
@@
The Hyper Suprime-Cam first-year data (HSC Y1), based on the intermediate data release S16a, covers $137\,\mathrm{deg}^{2}$ with an effective number density of $\sim 22\,\mathrm{arcmin}^{-2}$ (Mandelbaum2018_PASJ_70_25). Galaxies are assigned into $4$ tomographic bins based on photometric Ephor AB redshifts from $5$-band grizygrizy photometry (Hikage2019_PASJ_71_43). For our analysis, we make use of the S17A Wide-COSMOS catalog of the COSMOS field that is based on observations at the depth of the full survey, contains all objects passing the fiducial weak lensing selections cuts, and is matched against the COSMOS2015 catalog (Laigle2016_ApJS_224_24). In the default HSC first-year photometric redshift calibration, targets are weighted by the product of their lensing weights and a SOM-based weight to partially correct for cosmic variance in the COSMOS field. The fiducial $n(z)$ is then based on the weighted histogram of COSMOS2015 multi-band photometric redshifts, i.e., no spectroscopic redshifts were employed (Hikage2019_PASJ_71_43).
II.4 Kilo-Degree Survey
@@
The KiDS-1000 data release of the Kilo-Degree Survey covers $\sim 1000\,\mathrm{deg}^{2}$ with an effective number density of $6.2\,\mathrm{arcmin}^{-2}$ (Giblin2021_AA_645_105). A total of $5$ tomographic bins are assigned based on BPZ photometric redshifts from the $ugriZYJHK_{s}$ optical to near-infrared bands. For our analysis, we use a photometric catalog of the COSMOS field at a higher depth than the main survey that has been artificially degraded to simulate the depth of the KiDS-1000 weak lensing catalog (A. Wright, personal communication). Unlike for DES Y3 and HSC Y1, this KiDS-1000 COSMOS catalog has not been processed with the weak lensing pipeline. To estimate the impact of weak lensing cuts and weights, we also make use of the DR 4.1 photometric catalogs in combination with the KiDS-1000 weak lensing SOM-gold catalog, as described in section III.2. Finally, the fiducial photometric redshift calibration has been described in Hildebrandt2021_AA_647_124 and is based on a combination of SOM and clustering redshifts.
III Methods
@@ III.1 Incompleteness Correction
III.1.1 Fiber Assignment Probabilities
Ideally, the fiber assignment probability would be computed from first principles if the targeting criteria were well-defined and we had random realizations of the fiber assignment such as for the DESI large-scale structure catalogs (Lasker2025_JCAP_01_127). In practice, the DESI COSMOS-XMM catalog contains targets from a large heterogeneous set of secondary target programs, often without alternative realizations of fiber assignments.
@@ III.1.1 Fiber Assignment Probabilities
The values predicted by the trained network can be interpreted as the fiber fraction $p_{\mathrm{f}}$ given an object’s coordinates and photometry. To avoid spurious overfitting we employ a K-fold approach by always training the networks on $90\%$ of the data and calculating $p_{\mathrm{f}}$ for the remaining $10\%$, i.e., $K=10$. Finally, this process is repeated $100$ times for the whole sample such that each individual galaxy has $100$ predictions. We typically find the scatter between different neural network predictions $\sigma_{p_{\mathrm{f}}}$ for the same galaxy to be of order $5\%$ or less. Thus, after averaging those predictions, the scatter is of order $0.5\%$ or less. We note that while the network has access to the coordinates of objects, it is unlikely to faithfully reproduce small-scale fiber assignment effects such as fiber collisions and the influence of higher-priority targets (Lasker2025_JCAP_01_127). This is not a concern for our work since we do not study small-scale clustering. Instead, we are mainly concerned with having accurate mean fiber probabilities as a function of photometry within overlap regions with weak lensing surveys. Finally, we observe that the raw $p_{\mathrm{f}}$ do not reproduce the mean fiber rate $f_{\mathrm{f}}$ in $5\%$ bins of $p_{\mathrm{f}}$ perfectly and instead show small offsets. We mitigate this by manually adjusting the raw $p_{\mathrm{f}}$ values via $p_{\mathrm{f}}\rightarrow f_{\mathrm{f}}(p_{\mathrm{f}})$.
III.1.2 Importance Weights